Given utility maximization problem U= Q1Q2 subject to 10Q1 +2Q2=240 a. Derive the Lagrange function b. Derive the first order conditions c. Use Cramer’s rule to find the critical values of Q1, Q2 and �


1
Expert's answer
2021-04-16T07:38:59-0400



i.) Deriving the Lagrange function:

"Z = Q1Q2+ \u03bb(240-10Q1-2Q2)"

ii.) first-order conditions:

"ZQ1 = Q2− λ10 = 0\\


 ZQ2 = Q1− λ 2 = 0 \\


Zλ = 240 − 10Q1 −2 Q2 =0."


"Z\lambda=240-10Q1-2Q2=0"

"ZQ1=Q2- \lambda10=0"

"ZQ2=Q1-\lambda2=0"


iii.) "\begin{bmatrix}
 0 & -10 & -2 \\
 -10 & 0 &1 \\
-2 & 1 & 0
\end{bmatrix}" "\begin{bmatrix}
 \lambda \\
 Q1 \\ Q2
\end{bmatrix}" = "\begin{bmatrix}
 -240 \\
 0 \\ 0
\end{bmatrix}"


Q1M"=\frac{240}{2[-10]}= -12"


Q2M"=\frac{240}{2[-2]}=-60"


"\lambda=\frac{240}{2[-10.-2]}=6"